Hope someone can help explain this to me. I do not undertand.
A swimming pool holds540,000 liters of water. The pool has two drainage pipes. When the pool is completely full, the first pipe alone can empty it in 150 minutes, and the second pipe alone can empty it in 225 minutes. When both pipes are draining together, how long does it take them to empty the pool?

Thank you

Answer by nerdybill(7384)

A swimming pool holds540,000 liters of water. The pool has two drainage pipes. When the pool is completely full, the first pipe alone can empty it in 150 minutes, and the second pipe alone can empty it in 225 minutes. When both pipes are draining together, how long does it take them to empty the pool?
.
Let x = time to drain pool when both pipes are used
then
x(1/150 + 1/225) = 1
multiplying both sides by (150)(225):
x(225 + 150) = (150)(225)
225x + 150x = 33750
375x = 33750
x = 90 minutes